Given the following information for The Pelican Inn, what is the direct labor time variance?

Actual housekeeping staff hours worked 760
Standard staff hours for rooms cleaned 800
Standard staff pay rate per hour $12.50
Actual staff pay rate per hour $13.00

a.$500 unfavorable

b.$400 unfavorable

c.$(20) favorable

d.$(500) favorable
Answer #1

Labor time variance

= (SH-AH) *SR

= (800 - 760) * 12.50

= 40 * 12.50

= (500) Favourable

Option D is the answer